# Controlling Risk of Retrieval-augmented Generation: A Counterfactual Prompting Framework

### v1

@@ 1 Introduction
In this work, we propose a new task of risk control for RAG (RC-RAG) to decide whether to keep or discard the RAG outputs based on confidence assessment. We then introduce a novel counterfactual prompting framework for RAG under the zero-shot scenario, leveraging the counterfactual thinking for confidence assessment based on two latent factors. Counterfactual Pearl (2009) describes the human capacity to learn from prior experiences by imagining the outcomes of alternative actions that could have been taken. For a language model, we can inject counterfactual thinking into prompt, like “what if…” or “assume that”, to imagine or simulate the consequences of changing a factor. Here, we induce the model to imagine scenarios where the quality of the retrieved results and their usage are poor, then measure its confidence based on the effect of these imagined scenarios on the answers. Specifically, our framework consists of three major modules, i.e., a prompting generation module, a judgment module, and a fusion module: (i) the prompting generation module generates answers under two scenarios that challenges the improper use and poor quality of the retrieved results, respectively; (ii) the judgment module determines whether to discard or keep the generated answers for both scenarios; and (iii) the fusion module combines the judgment results from both scenarios to produce the final decision for selective output. In order to avoid overestimating confidence affecting risk control, the prompting generation module and the judgment module are executed iteratively to produce the judgment for each scenario. It is important to note that our method is a general post-processing technique, making it applicable to almost any existing RAG method.
@@ 2 Related work
Retrieval-augmented generation. The typical retrieval-augmented generation (RAG) method follows a retrieve-then-generate pipeline, first retrieving relevant documents from a grounding corpus and then generating the final answer by the frozen generators Shi et al. (2023); Ram et al. (2023). The retrieval augmentation is performed for all the questions through a single round Lewis et al. (2020); Guu et al. (2020); Izacard and Grave (2021); Shi et al. (2023) or multiple rounds Borgeaud et al. (2021); Ram et al. (2023); Trivedi et al. (2023); Jiang et al. (2023); Liu et al. (2024). However, such practice sometimes hurt generation performance, due to the unsatisfactory retrieved results Mallen et al. (2023); Ren et al. (2023); Yoran et al. (2023); Tan et al. (2024). The reason may lie in the inconsistency between the relevance judgments in retrieval stage and the utility judgments in generation stage Zhang et al. (2024). Besides jointly optimization of the retriever and generator Guu et al. (2020); Lewis et al. (2020); Singh et al. (2021); Izacard et al. (2023),
another solution is adaptive retrieval augmentation Jiang et al. (2023); Asai et al. (2023); Wang et al. (2023), which actively determines when to retrieve based on internal knowledge boundaries.
@@
Counterfactual thinking. As the third level of the causal ladder after association and intervention, counterfactual reflects causality by imagining “what would the outcome be had the variable(s) been different” Pearl (2009); Nan et al. (2021). Counterfactual inference helps model unchanging causal mechanisms for better generalization and debias, which can be utilized for text classification, visual question answering, recommendation system and so on Qian et al. (2021); Niu et al. (2021); Wei et al. (2021); Wang et al. (2022); Deng et al. (2023). It can calibrate causal effects through mediation analysis, by estimating the total effect and then eliminating the undesired effect Xie et al. (2021). Different from these works, we focus on injecting counterfactual thinking into the prompt to better apply RaLLMs.
@@ 4 Counterfactual prompting framework
In order to avoid overestimating confidence, we iteratively execute the prompting generation module and the judgment module in $N$ rounds to produce the judgment for each scenario. Specifically, if the answer remains the same after $N$ rounds, it can be judged as keep. Otherwise, if the answer changes once, it is judged as discard.
@@ 6.2 Impact of retriever
To answer RQ4, we compared the performance of risk control of RAG with different retriever. Results are shown in Table 4, conducted on the RC-NQ test set using Mistral as a generator.
@@ 7 Conclusion
Our approach requires multiple rounds of prompting generation, which is computationally expensive. Therefore, further exploration of more efficient prompting generation methods is necessary. Additionally, the current fusion strategy is heuristic. In the future, semantic information can be incorporated to aid in the integration of the two judgments. Furthermore, we only considered risk control in a zero-shot scenario. In the future, designing objective functions based on risk-related metrics for joint training with the RAG framework could be explored, aiming to achieve an effective balance between risk control and answering capability.

### v2

@@ 1 Introduction
In this work, we propose a new task of risk control for RAG (RC-RAG) to decide whether to keep or discard the RAG outputs based on confidence assessment. We then introduce a novel counterfactual prompting framework for RAG under the zero-shot scenario, leveraging the counterfactual thinking for confidence assessment based on two latent factors. Counterfactual Pearl (2009) describes the human capacity to learn from prior experiences by imagining the outcomes of alternative actions that could have been taken. For a language model, we can inject counterfactual thinking into prompt, like “what if…” or “assume that”, to imagine or simulate the consequences of changing a factor. Here, we induce the model to imagine scenarios where the quality of the retrieved results and their usage are poor, then measure its confidence based on the effect of these imagined scenarios on the answers. Specifically, our framework consists of three major modules, i.e., a prompting generation module, a judgment module, and a fusion module: (i) the prompting generation module generates answers under two scenarios that challenges the improper use and poor quality of the retrieved results, respectively; (ii) the judgment module determines whether to discard or keep the generated answers for both scenarios; and (iii) the fusion module combines the judgment results from both scenarios to produce the final decision for selective output. It is important to note that our method is a general post-processing technique, making it applicable to almost any existing RAG method.
@@ 2 Related work
Retrieval-augmented generation. The typical retrieval-augmented generation (RAG) method follows a retrieve-then-generate pipeline, first retrieving relevant documents from a grounding corpus and then generating the final answer by the frozen generators Shi et al. (2023); Ram et al. (2023). The retrieval augmentation is performed for all the questions through a single round Lewis et al. (2020); Guu et al. (2020); Izacard and Grave (2021); Shi et al. (2023) or multiple rounds Borgeaud et al. (2021); Ram et al. (2023); Trivedi et al. (2023); Jiang et al. (2023); Liu et al. (2024). However, such practice sometimes hurt generation performance, due to the unsatisfactory retrieved results Mallen et al. (2023); Ren et al. (2023); Yoran et al. (2023); Tan et al. (2024). The reason may lie in the inconsistency between the relevance judgments in retrieval stage and the utility judgments in generation stage Zhang et al. (2024). Besides jointly optimization of the retriever and generator Guu et al. (2020); Lewis et al. (2020); Singh et al. (2021); Izacard et al. (2023), another solution is adaptive retrieval augmentation Jiang et al. (2023); Asai et al. (2023); Wang et al. (2023), which actively determines when to retrieve based on internal knowledge boundaries.
@@
Counterfactual thinking. As the third level of the causal ladder after association and intervention, counterfactual reflects causality by imagining “what would the outcome be had the variable(s) been different” Pearl (2009); Nan et al. (2021). Counterfactual inference helps model unchanging causal mechanisms for better generalization and debias, which can be utilized for text classification, visual question answering, recommendation system and so on Qian et al. (2021); Niu et al. (2021); Wei et al. (2021); Wang et al. (2022); Deng et al. (2023). It can calibrate causal effects through mediation analysis, by estimating the total effect and then eliminating the undesired effect Xie et al. (2021). Different from these works, we focus on injecting counterfactual thinking into the prompt to better apply retrieval-augmented LLMs.
@@ 4 Counterfactual prompting framework
To reduce the likelihood of overestimating confidence, the prompting generation and judgment modules can be executed iteratively for $N$ rounds to validate the decision for each scenario. A decision is made as keep only if the answer remains consistent across all $N$ rounds. To balance computational efficiency, we have set $N$ to 1.
@@ 6.2 Impact of retriever
To answer RQ4, we compared the performance of risk control of RAG with different retrievers. Results are shown in Table 4, conducted on the RC-NQ test set using Mistral as a generator.
@@ 7 Conclusion
Firstly, the two latent factors influencing RAG’s confidence are human-defined, which may not encompass the full spectrum of risk sources. Future work could explore more diverse factors identified by LLMs, combined with statistic analysis.
Methodologically, our prompting generation approach is computationally intensive. Further exploration is needed to develop more efficient prompting strategies. The judgment module currently struggles with long answers, which requires a more sophisticated matching function. Additionally, the current fusion strategy is heuristic. Future enhancements could include semantic information to better integrate the two judgments.
Furthermore, we have focused solely on risk control in a zero-shot scenario. How to improve RAG answers in this scenario deserves further investigation. Also, designing objective functions based on risk-related metrics for joint training with the RAG framework could be explored, aiming for a balanced trade-off between risk control and response quality.